OpenStudy (jakeenglish):

Can someone please help me ASAP. It's urgent! Solve the following logarithmic equation. 2 log 3 - log x = 2 this is exactly how it's written, and i'd appreciate if someone explained the steps- not just told me the answer.

OpenStudy (anonymous):

Umm, does it not give a base? because then it isn't possible...

OpenStudy (jakeenglish):

it doesn't really, but i believe the base is supposed to be 3?

OpenStudy (jdoe0001):

so.... hmm

OpenStudy (jdoe0001):

\(\bf 2 log(3) - log(x) = 2\) is that it?

OpenStudy (jakeenglish):

yes i believe so! that's seriously how it's written though, not how it usually is with the bases very clear.

jhonyy9 (jhonyy9):

log(3^2 /x) = log2^2

OpenStudy (jdoe0001):

if no log base is given, "base 10" is the given

jhonyy9 (jhonyy9):

i think it so

jhonyy9 (jhonyy9):

@jdoe0001 - when i have learned it log was without base log base 2

OpenStudy (jdoe0001):

hmm have you got your log rules sheet around? check how to get the "2log(3)" simplified first off

OpenStudy (jdoe0001):

@jhonyy9 hmmm pretty sure is "10" when is just a bare "log"

OpenStudy (jdoe0001):

but one never knows, maybe it was teh book, or just the teacher

OpenStudy (jakeenglish):

i'm so confused, i believe 2log(3) is .954242509

jhonyy9 (jhonyy9):

but this i have learned log base ten just lg and without 10

OpenStudy (whpalmer4):

yes, \[\log(x) = \log_{10}(x)\]by convention

OpenStudy (jdoe0001):

@jakeenglish http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif <--- use the 3rd rule there, for the 1st term what does it give you?

jhonyy9 (jhonyy9):

so probably there in USA - but how i have learned it was log without base what mean log base 2 sorry

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ 2}}log(3)\qquad \qquad {\color{brown}{ a}}log(b)\iff log(b^{{\color{brown}{ a}}})\) so...what would 2log(3) give you then?

OpenStudy (jdoe0001):

notice, the rule 3 listed there also notice the rule 2, we'll need that one shortly

OpenStudy (jakeenglish):

sorry sorry i was working on another problem

OpenStudy (jdoe0001):

yes? no? maybe? no the foggiest? was busy chasing the dog?

OpenStudy (whpalmer4):

@jhonyy9 that's true in computer science, where most logs are binary, but in general, log(x) means either log base 10, or log base e (natural logarithm)

OpenStudy (jakeenglish):

so it would be \[\log(3^2)\] ?

OpenStudy (jdoe0001):

aha yes.... so... let's us use the 2nd rule listed now one that \(\bf log(3^2)-log(x)\qquad \qquad log_c(a)-log_c(b)\implies log_c\left( \frac{a}{b} \right)\) so.. what would that give you, using the 2nd rule on that sheet?

OpenStudy (jakeenglish):

give me one second, logs aren't my strong suit so sorry if this takes a little while for me to comprehend

OpenStudy (jdoe0001):

k

OpenStudy (jakeenglish):

i'm a little confused, because there's no base? or am i missing something?

jhonyy9 (jhonyy9):

@whpalmer4 - thank you - but i know what i have learned so lg mean log base ten and log natural was ln and not different way

OpenStudy (jakeenglish):

ok the site crashed for me but i think i figured it out a little bit \[\log(\frac{ 9 }{ x }) = 2\]

OpenStudy (joclyn16):

9 over 100 or .09

OpenStudy (jdoe0001):

,hmm yes

OpenStudy (jdoe0001):

hold the mayo

OpenStudy (anonymous):

log 3^3 - log x = 2 ==> log (3^3/x) = 2 ==> 3^3/x = e^2. ==> x = 3^3/e^2. Or if it's base 10 then x = 3^3/10^2. just apply the log rules.

OpenStudy (jdoe0001):

\(\bf log_{\color{red}{ a}}{\color{blue}{ b}}=y \iff {\color{red}{ a}}^y={\color{blue}{ b}}\qquad\qquad % exponential notation 2nd form {\color{red}{ a}}^y={\color{blue}{ b}}\iff log_{\color{red}{ a}}{\color{blue}{ b}}=y \\ \quad \\ \quad \\ \textit{now, notice what you have} \\ \quad \\ log\left( \cfrac{9}{x} \right)=2\implies log_{{\color{red}{ 10}}}\left( {\color{blue}{ \cfrac{9}{x} }} \right)=2\implies ?\) what does that change to?

OpenStudy (jdoe0001):

anyhow.. need to dash... so... ahemm notice \(\bf log_{\color{red}{ a}}{\color{blue}{ b}}=y \iff {\color{red}{ a}}^y={\color{blue}{ b}}\qquad\qquad % exponential notation 2nd form {\color{red}{ a}}^y={\color{blue}{ b}}\iff log_{\color{red}{ a}}{\color{blue}{ b}}=y \\ \quad \\ \quad \\ \textit{now, notice what you have} \\ \quad \\ log\left( \cfrac{9}{x} \right)=2\implies log_{{\color{red}{ 10}}}\left( {\color{blue}{ \cfrac{9}{x} }} \right)=2\implies {\color{red}{ 10}}^2={\color{blue}{ \cfrac{9}{x} }}\) pretty sure you can take it from there